OpenStudy (anonymous):

find the standard form of the equation of a line passing through the points B(0, 0), C(4, -4)

OpenStudy (popitree):

y=6

OpenStudy (anonymous):

thank you so much!

OpenStudy (popitree):

parallel to x axis cut y axis at 0,6

OpenStudy (anonymous):

thank you

OpenStudy (across):

\(y=6?\) How is that right?

OpenStudy (across):

The right answer is \(y+x=0\).

OpenStudy (popitree):

the qstn changed i think, the qstn was passing through 0.6 and 5,6

OpenStudy (popitree):

yes, y = -1.x + 0

OpenStudy (across):

I see.

OpenStudy (mrnood):

@Shelby91953 you should always be very careful of people giving answer without explanation - anndf CHECK IT yourself @popitree is incorrect for the currently posted question

OpenStudy (anonymous):

yeah I changed the question sorrry

OpenStudy (popitree):

i was right then :)

OpenStudy (popitree):

thanks for confession

OpenStudy (anonymous):

so the answer to the question NOW is y + x=0?

OpenStudy (popitree):

yes

OpenStudy (mrnood):

@Shelby91953 But - do you understand HOW to get that answer? (because the NEXT question of this topic will NOT have an answer y+x=0)

OpenStudy (anonymous):

no lol